Density is a physical property of a molecule that’s measured as mass per unit volume. Given that density is usually expressed in the units of g/cm3, calculate the density of a cube of plastic that has a mass of 2.331 g and an edge length of 1.500 cm (volume of a cube is V = l3). Be sure to show all steps in the process and include units throughout the problem.

1
Expert's answer
2022-03-22T17:49:02-0400

Mass of cube = 2.331 g


Edge length of cube = 1.500 cm


Volume of cube= "l^3"


Volume ="(1.500)^3" cm3

Volume = 1.500×1.500×1.500= 3.375 cm3


"Density=\frac{mass}{volume}= \frac{2.331}{3.375}=" 0.691 "\frac{g}{cm^3}"